OpenStudy (smile.89745):

Consider the following equations: −x − y = 1 y = x + 3 If the two equations are graphed, at what point do the lines representing the two equations intersect? (−1, 2) (−2, 1) (1, −2) (2, −1)

OpenStudy (smile.89745):

PLease help i tired all of them and the third one is not correct

OpenStudy (smile.89745):

@SkyVoltage @drawwithapurpose

OpenStudy (goldenticket):

The answer is A

OpenStudy (smile.89745):

Can you explain to me how @GoldenTicket

OpenStudy (goldenticket):

Yea do you got a graph to show?

OpenStudy (smile.89745):

And can you help me on some more

OpenStudy (smile.89745):

no sorry it did not come with one

OpenStudy (goldenticket):

Mhmm weird it was to show a graph oh well if you just draw a graph and graph the points to the right sector of the location and add them that is how you get your answer

OpenStudy (smile.89745):

and can you help me with these: Consider the following set of equations: Equation C: y = 2x + 8 Equation D: y = 2x + 2 Which of the following best describes the solution to the given set of equations? No solution One solution Two solutions Infinite solutions Consider the following pair of equations: y = 3x + 3 y = x − 1 Explain how you will solve the pair of equations by substitution. Show all the steps and write the solution in (x, y) form.

OpenStudy (smile.89745):

Can you walk me through the steps to if you want?

OpenStudy (smile.89745):

@SkyVoltage @dame @just_one_last_goodbye @MollyNicole2015 PLease i need help